Suppose that the demand and supply functions are P=10-Q and P=Q respectively. Which of the following statements are correct for the value of consumer and producer surplus in the equilibrium?

Select one or more:

a. The value of producer surplus is 25.

b. The value of consumer surplus is 15.

c. The value of consumer surplus is 12.5.

d. The value of producer surplus is 15.

e. The value of consumer surplus is 25.

f. The value of producer surplus is 12.5.
Answer #1

Demand function P = 10-Q

Supply function P =Q

Finding equilibrium point

Q = 10 - Q

Q = 5, At equilibrium price = 5

Finding intercept of demand function on y axis,

When Q = 0, P =10 (from demand function)

Finding intercept of supply function on y axis,

When Q = 0, P =0 (from supply function)

Consumer surplus = 0.5 * 5 * (10-5) = 0.5 * 5 * 5 = 12.5

Producer surplus = 0.5 * 5 * 5 = 12.5

So c & f are correct options